OpenStudy (anonymous):

a container contains 50 green tokens, 15 blue tokens, and 3 red tokens. two tokens are randomly selected without replacement compute P(FIE). E-you select a non green token first F-the second token is green P(FIE)=??

OpenStudy (anonymous):

P(F|E)=P(F intersection E)/P(E) P(E+F)= (18/68)*(50/67) P(E)=18/68 So P(F|E)=(18/68)*(50/67) /(18/68)= 50/67 P(F) would be 50/68, so P(F|E) is a bit more likely

OpenStudy (anonymous):

would that be reduced to 25/64

OpenStudy (anonymous):

No, Andras is saying the answer would be 50/67, with which I agree. :)

OpenStudy (anonymous):

how did you get 25/64?

OpenStudy (anonymous):

you wrote 50/68

OpenStudy (anonymous):

that is P(F) on its own, P(F|E)=50/67

OpenStudy (anonymous):

do you know what is the difference?

OpenStudy (anonymous):

so what would the P(FIE) be for this problem then a container contains 30 green tokens ,20 blue tokens, and 4 red tokens. two tokens are randomly selected without replacement compute P(FIE). E-you select a non red token first F- the select token is red

OpenStudy (anonymous):

You can do the work here :) I will double check it for you, is that fine?

OpenStudy (anonymous):

you need to count the probability that they occur together and the probability of P(E)